OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = –5 and a6 = –5,120?

OpenStudy (anonymous):

i cant seem to get the right answer, something is wrong in my steps

OpenStudy (kropot72):

The equation for the 6th term is: \[-5x ^{(6-1)}=-5120\] \[5x ^{5}=5120\] x = 4 Did you get this?

OpenStudy (anonymous):

i have to use the formula an=a1r ^n-1 i can get all the way through the steps its just when i get where you have to take the square root its looks like this \[\sqrt[7]{1024?}\] i dont know how to do that

OpenStudy (kropot72):

The 9th term is found as follows: \[-5\times 4^{8}=-327680\]

OpenStudy (anonymous):

how did you get -5 x 4^8 ? like where did yiu get it from ?

OpenStudy (kropot72):

Actually you needed to take the 5th root of 1024: \[\sqrt[5]{1024}=(1024)^{1/5}=4\] To calculate it I use the following function key on a calculator: \[y ^{x}\] entering 1024 for y and 1/5 for x

OpenStudy (anonymous):

but why 5 and not 7 ?

OpenStudy (kropot72):

Because the formula asks for (the position of the term - 1). So the calculation for the 6th term will have the power (6 - 1) = 5

OpenStudy (anonymous):

it as for the 9th twerm

OpenStudy (kropot72):

Good :) The power in the calculation for the 9th term is (9 - 1) = 8

OpenStudy (anonymous):

so i still dont get how to take the square root ? \[\sqrt[8]{1024?}\]

OpenStudy (kropot72):

The common ratio r = 4 So using the formula you were given to find the 9th term you do the following: (first term) * 4^(9 - 1) =\[-5\times 4^{(9-1)}=-5\times 4^{8}=-327680\]

OpenStudy (anonymous):

i get that, where are you gettinh the 4 from? since it didnt give you the common ratio

OpenStudy (kropot72):

In the first posting i used x = common ratio. Your formula calls it r. So x = r = 4 = common ratio

OpenStudy (anonymous):

Do you mind working out thr problem ? cause im very confused.

OpenStudy (kropot72):

I will get back to u in a few minutes :)

OpenStudy (anonymous):

okat thank you

OpenStudy (kropot72):

The formula looks like this. \[a _{n}=a _{1}\times r ^{(n-1)}\] a is a term in the series and n stands for the position. Suffix 1 after a identifies it as the first term. r is the common ratio To find r (common ratio) substitute the value that is given for the 6th term into the formula as follows: \[-5120=-5\times r ^{(6-1)}\] \[5120=5\times r ^{(6-1)}=5\times r ^{5}\] \[r ^{5}=5120\div5=1024\] \[r=\sqrt[5]{1024}=4\] Can you follow?

OpenStudy (anonymous):

oh yes, i see you to the 6 from a6

OpenStudy (kropot72):

\[a _{6}\]is the 6th term

OpenStudy (anonymous):

yes then 6-1 is how you got 5, and yiu turn that into 1/5 ?

OpenStudy (kropot72):

\[\sqrt[5]{1024}=1024^{(1/5)}\]

OpenStudy (anonymous):

okay thank you